# Joint Liability Model with Adaptation to Climate Change

### v1

@@ 3.4 Joint Liability Contracts with a Group Size of $n$
In this sub-section, we present a model which formalizes the assumption that the group of borrowers is collectively responsible for the entire group loan (i.e., $nw$). The expected utility of a borrower $i$ who adopts a repayment strategy during any period when they receive financing is specified as
@@
Constraints similar to those in the case of a group size of two are as follows:
@@
    There is at least one borrower who is successful and capable of paying the entire group’s repayment amount.
@@
    The lender must be able to sustain the lending game over the periods or at least break even. Thus, the expected repayment amount of each borrower has to be at least as large as $L(1+\varepsilon)$, where $\varepsilon$ is the risk-free rate. Still, given that all constraints have been established under the presumption that the group will successfully repay the total loans, in order to ensure their eligibility for subsequent loan rounds, we have not factored in the scenario where all borrowers within the group experience failure on the LHS of this constraint.
@@
    Similarly, we can get $w\geqslant\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$. Thus, the value of $w$ should be set as $\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$ due to the negative relationship between $w$ and $\pi_{i}$.
After substituting $w=\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$ in the objective function (Equation 3.20), the ex-ante expected payoff function for the borrower $i$ in a size $n$ joint liability group is represented as
@@
Proposition 4: The optimal individual-ESE score and group size $n$ have a negative relationship.
A higher ESE score indicates a lower default risk, which means that the borrower is more creditworthy and can obtain a loan with a smaller group size. On the other hand, each member of the group is jointly responsible for the repayment of the loan, and if one member defaults, the other members will have to repay the outstanding amount. Thus, a low-quality borrower may choose to loan with a larger group to reduce the probability of default and minimize the potential burden on any one individual. Next, we verify this intuition based on the following derivations.
Take the first-order partial derivative of the expected payoff w.r.t $E$ to get the optimal individual-ESE score for a borrower:
@@
Since it is hard to find an explicitly analytical form of $n$ and $E$, we adopt an Implicit function derivation. Let $E(n)$ replace $E$ and take the partial derivative w.r.t. $n$ on both sides of $kp\overline{Y}+p\underline{Y}\left[n(1-kE-b)^{n-1}-1\right]-kc(kE+b)=0$. Then we can get the following result:
@@
After reorganizing the formula, we get the partial derivative of $E(n)$ w.r.t $n$ as
@@
Obviously, the denominator in Equation 3.28 is greater than 0. When $n\geqslant 1$ and $e\in(0,1)$, then $n\log(1-e)<-1$, and $(1-e)^{n-1}[1+n\cdot\log(1-e)]<0$. Thus, the numerator is less than 0. As a result, $\frac{\partial E(n)}{\partial n}<0$, which proves the optimal individual-ESE score and group size $n$ have a negative relationship.
In light of the negative correlation between optimal individual-ESE scores and group size, it is recommended that insurance companies prioritize lenders with higher ESE scores when the group size is smaller to mitigate potential risks. Borrowers with high scores are typically perceived to exhibit positive economic, environmental, and social performance, which leads to higher expectations for sustained success and repayment probability over the long term. Despite the relatively small group size, lenders may still have heightened confidence in the ability of high-scoring borrowers to demonstrate honesty and success, and to exhibit willingness to repay their loans. As a result, lenders may consider reducing group size requirements for high-scoring borrower groups. Conversely, as group size increases, the requirement for lenders’ ESE scores may potentially decrease, as a larger group may possess a greater risk mitigation capacity and an increased likelihood of successfully repaying the group loan. Consequently, lenders may opt to reduce the score requirements in such scenarios.
Proposition 5: The optimal individual-ESE score converges to a constant level when the group size reaches a large number.
The Equation 3.28 tells us that $\lim_{n\rightarrow\infty}\frac{\partial E(n)}{\partial n}=0$, which means that the optimal individual-ESE score converges to a constant level when the group size reaches a certain large number. To find the limit optimal score value, we still need to recall the first-order condition of the ex-ante expected payoff, which is
@@
Similarly, we set $C(e)=\frac{1}{2}ce^{2}$, then
@@
When $n\rightarrow\infty$, $n(1-kE-b)^{n-1}\rightarrow 0$, then Equation 3.29 can approach to $kp(\overline{Y}-\underline{Y})-kc\cdot(kE+b)$. Let $kp(\overline{Y}-\underline{Y})-kc\cdot(kE+b)=0$, the limit of the optimal ESE score will be expressed as a constant:
@@
To illustrate this proposition, we provide a numerical example, which is shown in Figure 2. Set the unit price $p=1$, the successful probability $e=k\times E+b=1/100\times E+0$, high yield amount $\overline{Y}=1000$, low yield amount $\underline{Y}=500$, and perform a simulation under the different group size $n\in[1,100]$.
@@
In situations where the group size is small, insurance companies may seek to attract borrowers with high ESE scores in order to mitigate risk. However, as the group size expands, the marginal benefit of an individual effort on group outcomes diminishes, resulting in a reduction of an individual farmer’s optimal effort level. This can create incentives for farmers to contribute less effort, leading to a ‘free-rider’ problem and decreased profits for the group. Therefore, the threshold value of the optimal ESE score can serve as a reference point for both lenders and borrowers, enabling them to avoid an inefficient resource allocation and a free-rider problem.
@@
4 Joint Liability Contracts under a Mean-Variance Utility Framework
4.1 Crop yields under climate change
Climate change refers to the long-term alterations in the Earth’s climate system, including changes in temperature, precipitation patterns, and sea levels, which are primarily caused by human activities. These changes are having major impacts on the agriculture sector of the economy, which is heavily dependent on weather patterns, soil fertility, water resources, and pest and disease pressures. Weather patterns are becoming more extreme and unpredictable due to climate change. This can cause droughts, floods, and heat waves, which can all negatively impact crop yields. For example, droughts can cause soil to become dry and hard, reducing plant growth and leading to low yields, while floods can wash away topsoil and damage crops. Extreme heat can also reduce yields by causing plants to wilt or preventing pollination. Soil fertility is also being affected by climate change. Higher temperatures and changes in rainfall patterns can lead to soil erosion, nutrient depletion, and soil compaction. Soil erosion can remove valuable topsoil, which is necessary for plant growth, while nutrient depletion can reduce the number of nutrients available to plants, leading to lower yields. Soil compaction can make it difficult for plant roots to penetrate the soil, reducing their ability to absorb water and nutrients. Water resources are also being impacted by climate change. Changes in precipitation patterns can cause droughts or floods, which can reduce water availability for crops. In addition, warmer temperatures can increase water demand, causing crops to require more water for growth. This can put a strain on already limited water resources, leading to competition between different water users, including farmers. Pest and disease pressures are also being affected by climate change. Changes in temperature and precipitation can alter the behavior of pests and disease-causing organisms, leading to outbreaks or changes in distribution. This can lead to increased crop damage and lower yields.
The Actuaries Climate Index (ACI) is an innovative climate monitoring tool that provides a comprehensive assessment of changes in extreme weather conditions and sea level rise in the United States and Canada. The ACI is a joint initiative of the American Academy of Actuaries and the Canadian Institute of Actuaries, designed to provide policymakers, business leaders, and other stakeholders with a reliable source of information on climate trends and their potential impacts. The ACI is based on the analysis of six climate indicators, including high and low temperatures, heavy precipitation, drought, and sea level rise. The index measures changes in these indicators over time and across regions, providing a clear picture of how the climate is changing in different parts of North America. One potential application of the ACI is to use it as a tool for predicting agricultural yields. Climate change can have significant impacts on agriculture, and the ACI can provide valuable information on how extreme weather events and other climate factors are likely to affect crop yields in different regions. Pan et al. (2022) used regression analysis to investigate the relationship between the ACI and crop yield and found that the ACI is a significant predictor of crop yield and (re)insurance ratemaking. Jiang and Weng (2019) examine the impact of climate change risk on agriculture-related stocks in Canada and the United States. To measure climate change risk, they use the ACI as a proxy and find that ACI trends have an adverse effect on agricultural production. They then assess the correlation between ACI trends and corporate profitability of agriculture-related companies, finding that companies in regions with higher climate change risk have worse financial performance. Based on these findings, they construct a risk-adjusted stock trading strategy that adjusts to climate change risk and find that the ACI is predictable on stock returns, indicating market inefficiency towards climate change risk. Their study contributes to the literature by using ACI as a reliable proxy for climate change risk and providing evidence of the mispricing of risk in the stock market.
4.2 The optimal ESE score with different risk aversion parameters
@@
When borrowers consider potential projects, they want to know not only what the most likely outcome is, but also how likely it is that they will earn a high return or experience a significant loss. The variability of the profit, or the range of possible returns, is an important consideration. If the potential variability of returns is too high or the risk of significant losses is too large, the borrower may decide not to invest in the venture. As far as I am aware, the current literature on joint liability problems only focuses on the expected return or expected profit. This means that they concentrate solely on the central point of an individual’s profit without taking into account the distribution of potential profits or the variability of these returns. However, focusing only on the expected return can be limiting, as it does not account for the risks. Therefore, we following incorporate the risk (variance) associated with expected profit into our utility function and conduct further analysis of mean-variance utility in the context of joint liability problems. The Mean-Variance Utility function for an individual farmer can be incorporated in his/her ex-ante expected profit function in Equation 4.1, where $P$ (a random variable) represents the expected profit and $\gamma$ represents the risk aversion parameter of this farmer.
@@
In order to derive the required optimal individual-ESE score, it is needed to take the first-order partial derivative of the ex-ante profit w.r.t the ESE score.
@@
It is hard to derive the analytical solution of $E$ score; so we instead use simulation studies to find some insightful patterns.
@@
Recall that the probability of achieving success in a project is mathematically represented as $e=k\times ESE+b$, wherein the parameters $k$ and $b$ are predetermined. Subsequently, the probability may be disaggregated into two constituent components. Specifically, the product of $k$ and $ESE$ represents the contribution of personal impact towards project success, whereas $b$ signifies the constant contribution of exogenous factors beyond individual control. The value of $k$ is utilized to calibrate the impact of $ESE$ on project success, while $b$ pertains to the influence of exogenous factors, including climatic and geographical conditions. To determine the value of $b$, the Actuaries Climate Index (ACI) for each region can be utilized as a reference. It is commonly understood that favorable climatic conditions have a positive effect on crop yields, which, in turn, increases the probability of project success. Conversely, unfavorable climatic conditions have an adverse impact on the likelihood of success. Hence, we delve deeper into the impact of varying climatic conditions caused by climate change on ESE score requirements. By changing the farmer’s risk aversion from 0 to 1 under different cost parameters, a series of simulations are conducted for a high success probability project under good climatic conditions (where $b=0.7$, see in Figure 3), a neutral probability project under neutral climatic conditions (where $b=0.5$, see in Figure 4), and a relatively low probability project under bad climatic conditions (where $b=0.3$, see in Figure 5). After observing and comparing these three figures, we conclude our analysis with three findings below. In general, any factor that can increase the success probability of the project and the repayment probability can directly or indirectly reduce the requirements of the ESE score.
  1. 1.
    By examining Figure 3, Figure 4, and Figure 5, we notice that as the probability of success decreases, the ESE score requirement continues to increase if the control cost remains constant. That is, the probability of a project’s success is inversely proportional to the required ESE score. This is because a higher success rate leads to an increased farmers’ ability to repay their loans, thus lowering the risk of default for the insurance company. As a result, the insurance company can lower its ESE score requirements when lending to farmers.
  2. 2.
    If we focus on any one of the three following figures, it is easy to find that as the risk aversion increases, the required ESE score will decrease monotonically. That is because an increase in farmers’ risk aversion score leads to more cautious contract behavior, reducing the probability of strategic default and increasing the probability of repayment. In other words, a higher degree of risk aversion among farmers implies a lower level of risk for the insurance company, leading to a lower ESE score requirement.
  3. 3.
    Still considering any one of the three figures, an increase in the cost parameter will result in a higher requirement for the ESE score. An increase in the cost parameter raises the production expenditure of farmers and burdens their economic resources, thus reducing their profits. Therefore, to avoid the high default probability associated with high costs, the corresponding ESE score requirement should increase. Hence, as the cost parameter increases, the insurance company should raise its ESE score requirement to reduce the risk of default among farmers.
Figure 3: High Successful Probability (Good Climatic Conditions)
Figure 4: Neutral Successful Probability (Neutral Climatic Conditions)
Figure 5: Low Successful Probability (Bad Climatic Conditions)
@@
Finally, we are investigating how changes in crop yields can impact the required score for ESE score. To do this, we vary the values of yields and observe the effect on the required ESE score. Specifically, we consider two scenarios: a ‘Low Yield Case’ where $\overline{Y}=1,000$ and $\underline{Y}=500$, and a ‘High Yield Case’ where $\overline{Y}=600$ and $\underline{Y}=300".$ The results of our simulation in Figure 6 show that there is a negative correlation between crop yields and the required ESE score, meaning that as crop yields increase, the required ESE score decreases. Conversely, as crop yields decrease, the required ESE score increases. When a borrower has lower yields, this can be an indication of lower income and potentially higher default risk. In other words, the higher requirement for borrowers with lower yields is a reflection of the lender’s risk management strategy. By setting stricter requirements for borrowers with lower yields, the lender is attempting to minimize their exposure to financial risk and ensure that borrowers are able to repay their loans on time and in full.

### v2

@@ 2.4 Joint Liability Contracts with a Group Size of $n$
In this subsection we present an individual ESE–joint liability model which formalizes the assumption that the group of borrowers is collectively responsible for the entire group loan (i.e., $nw$). The expected utility of a borrower $i$ who adopts a repayment strategy during any period when they receive financing is specified as
@@
Constraints similar to those in the case of a group size of two are as follows.
@@
    There is at least one borrower who is successful at his/her project and capable of paying the entire group’s repayment amount.
@@
    The lender must be able to sustain the lending game, or at least break even, over the periods. Thus the expected repayment amount of each borrower has to be at least as large as $L(1+\varepsilon)$, where $\varepsilon$ is the risk-free rate. It is again important to note that, given that all constraints have been established under the presumption that the group will successfully repay the total loans in order to ensure their eligibility for subsequent loan rounds, we have not factored in the scenario where all borrowers within the group experience failure on the left hand side of this constraint.
@@
    Similarly, we obtain $w\geqslant\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$. Thus the value of $w$ should be set as $\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$ due to the negative relationship between $w$ and $\pi_{i}$.
After substituting $w=\frac{L(1+\varepsilon)}{1-(1-e)^{n}}$ in the objective function (which is Equation (2.20)), the ex-ante expected payoff function for the borrower $i$ in a size $n$ joint liability group is represented as
@@
PROPOSITION 4 (Optimal individual–ESE score and group size $n$ have a negative relationship). A higher ESE score indicates a lower default risk, which means that the borrower is more creditworthy and can obtain a loan with a smaller group size. On the other hand each member of the group is jointly responsible for the repayment of the loan, and if one member defaults, the other members will have to repay the outstanding amount. Thus a low-quality borrower may choose to loan with a larger group to reduce the probability of default and minimize the potential burden on any one individual. Next we verify this intuition based on the following derivations. We take a first-order partial derivative of the expected payoff w.r.t $E$ to obtain an optimal individual–ESE score for a borrower
@@
Since it is hard to find an explicitly analytical form of $n$ and $E$, we adopt an implicit function derivation below. Let $E(n)$ replace $E$ and take the partial derivative w.r.t. $n$ on both sides of $kp\overline{Y}+p\underline{Y}\left[n(1-kE-b)^{n-1}-1\right]-kc(kE+b)=0$. Then we obtain the following result
@@
After reorganizing the formula, we can express the partial derivative of $E(n)$ w.r.t $n$ as
@@
It is clear that the denominator in Equation (2.28) is greater than 0. When $n\geqslant 1$ and $e\in(0,1)$, then $n\log(1-e)<-1$, and $(1-e)^{n-1}[1+n\cdot\log(1-e)]<0$. Thus, the numerator is less than 0. As a result $\frac{\partial E(n)}{\partial n}<0$, which proves the optimal individual–ESE score and group size $n$ have a negative relationship. In the light of the negative correlation between optimal individual–ESE scores and a group size, it is recommended that insurance companies prioritize lenders with higher ESE scores when the group size is smaller to mitigate potential risks. Borrowers with high scores are typically perceived to exhibit positive ESE performance, which leads to higher expectations for sustained success and repayment probability over the long term. Despite the relatively small group size, lenders may still have heightened confidence in the ability of high–scoring borrowers to demonstrate honesty and success, and to exhibit willingness to repay their loans. As a result lenders may consider reducing group size requirements for high-scoring borrower groups. Conversely, as the group size increases, the requirement for lenders’ ESE scores may potentially decrease, as a larger group may possess a greater risk mitigation capacity and an increased likelihood of successfully repaying the group loan. Consequently lenders may opt to reduce the score requirements in such cases. PROPOSITION5 (Optimal individual–ESE score converges to a constant level for a large enough group size). Equation (2.28) tells us that $\lim_{n\rightarrow\infty}\frac{\partial E(n)}{\partial n}=0$, which means that the optimal individual–ESE score converges to a constant level when the group size reaches a certain large enough number. To find the limit of the optimal score value, we recall the first-order condition of the ex-ante expected payoff, which is
@@
Similarly we set $C(e)=\frac{1}{2}ce^{2}$, then
@@
When $n\rightarrow\infty$, $n(1-kE-b)^{n-1}\rightarrow 0$, then Equation (2.29) approaches $kp(\overline{Y}-\underline{Y})-kc\cdot(kE+b)$. Let $kp(\overline{Y}-\underline{Y})-kc\cdot(kE+b)=0$, the limit of the optimal ESE score will be expressed as a constant
@@
To illustrate this proposition, we provide a numerical example, which is shown in Figure 2. We set the unit price $p=1$, the successful probability $e=k\times E+b=1/100\times E+0$, high yield amount $\overline{Y}=1000$, low yield amount $\underline{Y}=500$, and perform a simulation under the different group size $n\in[1,100]$.
@@
In situations where the group size is small, insurance companies may seek to attract borrowers with high ESE scores in order to mitigate risk. However, as the group size expands, the marginal benefit of an individual effort on group outcomes diminishes, resulting in a reduction of an individual farmer’s optimal effort level. This can create incentives for farmers to contribute less effort, leading to a ‘free-rider’ problem and decreased profits for the group. Therefore the threshold value of the optimal $ESE$ score can serve as a reference point for both the lenders and borrowers, enabling them to avoid an sub-optimal resource allocation and a free-rider problem.
@@
3 Joint Liability Contracts under a Mean–Variance Utility Framework
3.1 The optimal $ESE$ score with different levels of risk aversion
@@
When borrowers consider potential projects, they would like to know how likely it is that they will be able to earn a high return or experience a substantial loss. The variability of profits, or the range of possible returns, is an important consideration for these borrowers. If the potential variability of returns is too high or the risk of substantial losses is too large, the borrower may decide not to invest in the venture. To the best of our knowledge, the current literature on joint liability models focuses only on the expected return or expected profit. This means that they concentrate solely on the central point of an individual’s profit without taking into account the distribution of potential profits or the variability of these returns. However, focusing only on the expected return can be limiting, as it does not account for the risks. In this paper we incorporate the risk (proxied by variance) associated with expected profit into our utility function and conduct further analysis of mean–variance utility in the context of the individual ESE–joint liability model. Specifically the mean–variance utility function for an individual farmer can be incorporated in his/her ex–ante expected profit function in Equation (3.1), where $P$ (a random variable) represents the expected profit and $\gamma$ represents a risk aversion parameter of the farmer
@@
In order to derive the required optimal individual–ESE score in this setting, we take a first-order partial derivative of the ex-ante profit with respect to the ESE score
@@
Unfortunately the resulting analytical solution of the ESS score involves a rather complex expression and is difficult to interpret intuitively; so we resort to simulation studies to reveal several insightful patterns. Recall that the probability of achieving a success in a project is given by $e=k\times ESE+b$, wherein the parameters $k$ and $b$ are predetermined. Subsequently the probability may be disaggregated into two constituent components. Specifically the product of $k$ and $ESE$ represents a contribution of a personal impact towards a project success, whereas $b$ represents a constant contribution of exogenous factors beyond the individual control. The value of $k$ is utilized to calibrate the impact of $ESE$ on a project success, while $b$ pertains to the influence of exogenous factors, including climatic and geographical conditions. To determine the value of $b$, an Actuaries Climate Index (ACI) for each region can in principle be utilized as a reference. Briefly the ACI is based on an analysis of six climate indicators, including high and low temperatures, heavy precipitation, drought, and sea level rise. The index measures changes in these indicators over time and across regions, providing indication on how the climate is changing in different parts of a region. It is commonly understood that favorable climate conditions have a positive effect on crop yields, which, in turn, increases the probability of project success. Conversely, unfavorable climate conditions have an adverse impact on the likelihood of success. Hence, we delve deeper into the impact of varying climate conditions caused by climate change on ESE score requirements. By changing the farmer’s degree of risk aversion from 0 to 1 under different cost parameters, a series of simulations are conducted for a high success probability project under favorable climate conditions (where $b=0.7$ in Figure 3), a neutral probability project under neutral climate conditions (where $b=0.5$ in Figure 4), and a relatively low probability project under unfavorable climate conditions (where $b=0.3$ in Figure 5). After observing and comparing these three figures, we conclude our analysis with three findings below. In general any factor that can increase the success probability of the project and the repayment probability can directly or indirectly reduce the requirements of the $ESE$ score. By examining Figures 3, 4, and 5, we notice that, as the probability of success decreases, the $ESE$ score requirement continues to increase if the control cost remains constant; i.e., the probability of a project’s success is inversely proportional to the required ESE score. This is because a higher success rate leads to an increased farmers’ ability to repay their loans, thus lowering the risk of default for the insurance company. As a result the insurance company can lower its $ESE$ score requirements when lending to farmers. From inspecting any one of the three following figures, we see that, as the degree of risk aversion increases, the required ESE score decreases monotonically. This is because an increase in farmers’ risk aversion score leads to a more cautious contract behavior, reducing the probability of a strategic default and increasing the probability of repayment. In other words, a higher degree of risk aversion on the part of the farmers implies a lower level of risk for the insurance company, leading to a lower ESE score requirement. Again, from inspecting any one of the three figures, we can conclude that an increase in the cost parameter will result in a higher requirement for the ESE score. An increase in the cost parameter raises the production expenditure of farmers and burdens their economic resources, thus reducing their profits. Therefore, to avoid a high default probability associated with high costs, the corresponding ESE score requirement should increase. Hence, as the cost parameter increases, the insurance company should raise its ESE score requirement to reduce the risk of default among farmers.
Figure 3: High Successful Probability (Favorable Climate Conditions)
Figure 4: Neutral Successful Probability (Neutral Climate Conditions)
Figure 5: Low Successful Probability (Unfavorable Climate Conditions)
@@
Finally we arrive at the question of how changes in crop yields can impact the required $ESE$ score. To answer this question, we vary the values of yields and observe the effect on the required ESE score. Specifically we consider two scenarios: a ‘Low Yield Case’ where $\overline{Y}=1,000$ and $\underline{Y}=500$, and a ‘High Yield Case’ where $\overline{Y}=600$ and $\underline{Y}=300".$ The results of our simulation in Figure 6 show that there is a negative correlation between crop yields and the required ESE score, i.e., as crop yields increase, the required ESE score decreases. Conversely, as crop yields decrease, the required ESE score increases. When a borrower has lower yields, this can be taken as an indication of a lower income and a potentially higher risk of default; in other words, a higher requirement for borrowers with lower yields is a reflection of the lender’s risk management strategy. By setting stricter requirements for the borrowers with lower yields, the lender is attempting to minimize their exposure to financial risk and ensure that the borrowers are able to repay their loans on time and in full.